# DeepVigor+: Scalable and Accurate Semi-Analytical Fault Resilience Analysis for Deep Neural Networks

### v1

@@ Abstract
Growing exploitation of Machine Learning (ML) in safety-critical applications necessitates rigorous safety analysis. Hardware reliability assessment is a major concern with respect to measuring the level of safety. Quantifying the reliability of emerging ML models, including Deep Neural Networks (DNNs), is highly complex due to their enormous size in terms of the number of parameters and computations. Conventionally, Fault Injection (FI) is applied to perform a reliability measurement. However, performing FI on modern-day DNNs is prohibitively time-consuming if an acceptable confidence level is to be achieved. In order to speed up FI for large DNNs, statistical FI has been proposed. However, the run-time for the large DNN models is still considerably long.
In this work, we introduce DeepVigor+, a scalable, fast and accurate semi-analytical method as an efficient alternative for reliability measurement in DNNs. DeepVigor+ implements a fault propagation analysis model and attempts to acquire Vulnerability Factors (VFs) as reliability metrics in an optimal way. The results indicate that DeepVigor+ obtains VFs for DNN models with an error less than 1% and 14.9 up to 26.9 times fewer simulations than the best-known state-of-the-art statistical FI enabling an accurate reliability analysis for emerging DNNs within a few minutes.
@@ I Introduction
In recent years, the rapid advancement of Artificial Intelligence (AI), in particular of Machine Learning (ML), has provided a unique opportunity for automation leading to AI exploitation in safety-critical applications [1, 2]. Deep Neural Networks (DNNs) are dominantly employed in ML for complex tasks such as image classification and object detection that have applications in the safety-critical domain [3, 4]. According to recent regulations such as the European Union’s AI ACT [5] and the US National AI Initiative Act [6], AI systems that may cause adverse impacts on people’s safety are considered high-risk and must be rigorously assessed before deployment. This reflects the importance of early-stage safety risk assessment for AI deployment.
One of the major concerns in safety-critical applications is the hardware reliability of systems, which are threatened by hardware faults and errors [7]. With the technology miniaturization, the rate at which hardware faults occur is continuously rising in logic and memory circuits since the sensitivity of transistors increases against soft errors, temperature variations, etc. [8, 9, 10]. As DNNs are penetrating such applications e.g., perception in automotive [11], their hardware reliability assessment is necessary during the design and development phase, not only for risk certification but also for eliminating reliability weaknesses as early as possible.
On the other hand, the size of DNNs, in particular Convolutional Neural Networks (CNNs), in terms of memory footprint and number of operations is rapidly growing [12, 13]. Fig. 1 presents the growing model size in terms of the number of MAC operations and parameters for emerging DNNs and Vision Transformers (ViT), leading to their higher accuracy [14]. This trend necessitates leveraging more complex and performant computing systems such as GPUs, TPUs, etc. [15]. The fact that these complex systems are prone to hardware faults [16, 17, 18] leads to a necessary and complex task of their hardware reliability assessment prior to deployment in safety-critical applications.
@@
Hardware faults result in the alteration of values at the application level and might lead to erroneous outputs in a digital system. Hardware faults could lead to catastrophic consequences when exploiting hardware DNN accelerators in safety-critical applications. As an example, a hardware fault may result in misclassifying a red traffic light at an intersection in an autonomous vehicle and jeopardize the life of its passengers, as shown in Fig. 2, generated by AI.
Figure 2: An example of hardware faults effect on an autonomous vehicle, generated by https://getimg.ai.
In this regard, several research works have been carried out to assess and enhance DNNs’ hardware reliability against hardware faults listed in the related surveys [19, 16, 20, 21, 22]. Hardware reliability of DNN accelerators is the likelihood of a DNN accelerator performing as it is trained during the deployment [23]. The source of reliability threats is hardware-originated (soft errors, electromagnetic, temperature variation, aging, etc.) influencing a DNN’s functionality by changing its parameters (e.g., weights) or activation values during run-time.
@@
To ensure a reliable DNN deployment, the first step is to extensively evaluate the performance and functionality of pre-trained DNNs against hardware faults. To achieve this objective, three main approaches are identified throughout the state-of-the-art, in a prior systematic literature review [22]: 1) Fault Injection (FI) methods in which faults are injected into the target system and simulated, 2) Analytical methods where faults effects are mathematically analyzed, and 3) Hybrid methods that combine FI and analytical methods.
@@
It was observed that the majority of existing research works adopt FI (based on simulation, emulation, or irradiation). In FI experiments, faults can be accurately modeled, and the behavior of a system is evaluated in the presence of faults. However, given the growing size of emerging DNNs and their accelerators [15, 24], obtaining a fully precise evaluation is unfeasible and impractical.
@@
Throughout the literature, multiple research works are presented to significantly reduce the simulation complexity of FI experiments for DNNs while maintaining statistical accuracy. It includes software simulation [25, 26, 27, 28], or hardware emulation [29, 30, 31]. Nonetheless, any FI experiment requires a certain level of statistical confidence, which is obtained by a considerably high number of repetitions and a multitude of fault locations [26]. Therefore, FI-based approaches for reliability assessment are inherently unscalable and take days and weeks using powerful computing resources such as GPU.
@@
On the other hand, analytical approaches are presented to tackle the scalability issue. Nonetheless, they cannot provide reliability-oriented metrics including Vulnerability Factors (VFs) which is an essential metric for reliability evaluation. Moreover, the accuracy of analytical methods is not comparable to that of FI-based methods. To tackle the existing issues in the literature, a prior semi-analytical method named DeepVigor [23] is proposed as an alternative to simulation-based FI. DeepVigor provides accurate VF for bits, neurons, and layers of CNNs demonstrated to be faster than FI by utilizing a fault propagation analytical method. However, it analyzes all neurons in a CNN, implying a huge search space to obtain vulnerability values, leading to a scalability issue for large and deep emerging CNNs.
@@
In this paper, we propose a scalable, fast, and accurate alternative to FI named DeepVigor+, which provides VF for layers and models of DNNs. DeepVigor+ takes advantage of an optimal fault propagation analysis within neurons and entire DNNs to derive VFs in an optimized way. To the best of our knowledge, DeepVigor+ is the first scalable semi-analytical alternative to FI with a comparable accuracy for resilience analysis of DNNs in the literature. The contributions of this paper are as follows:
@@
  • Introducing a scalable, fast and accurate resilience analysis method for DNNs called DeepVigor+ for deriving VFs for DNNs’ layers and models by analyzing both parameters and activations as well as by exploiting an optimal error propagation analysis. DeepVigor+ conducts a novel statistical approach based on stratified sampling unleashing a fast and accurate resilience analysis for large and deep emerging DNNs,
@@
  • Extensively demonstrating the effectiveness of DeepVigor+ by examining its error compared to FI, leading to meeting a 1% mean absolute error in obtaining VF. Furthermore, its scalability is evidenced against statistical FI, resulting in 14.9 to 26.9 times fewer simulations than a cutting-edge state-of-the-art statistical FI approach. Derived VF by DeepVigor+ can provide the possibility of reliability comparison and visualization between and within various DNN models in a few minutes for deep and large DNNs.
@@
DeepVigor+ is presented as an open-source tool (https://github.com/mhahmadilivany/DeepVigor), unlocking reliability analysis for emerging DNNs and enabling researchers to quickly assess their reliability and design and develop fault-tolerant DNNs.
@@
Table I presents the abbreviations frequently used in the manuscript. In the rest of the paper, Section II highlights the gap by overviewing the related works on statistical FI as well as analytical methods. Section III presents the methodology of DeepVigor+ leading to fast and accurate reliability analysis for CNNs. Section IV explains how DeepVigor+ is implemented and describes the experiments, and their results are presented in detail in Section V to demonstrate the efficiency of DeepVigor+. Finally, Section VI concludes the paper.
@@ II Related Works: Resilience Analysis Methods for DNNs
With the fast growth of emerging DNNs’ size regarding their memory and computation requirements, their reliability analysis becomes prohitively complex, time-consuming and resource-hungry. In a recent literature review [22], it has been observed that the majority of papers studying the reliability of DNNs exploit a Fault Injection (FI) approach. Software-level simulation FI is widely adopted in reliability analysis due to its low design time and fast execution.
Multiple open-source tools for software-level simulation-based reliability analysis are proposed in the literature, and the most adopted ones include PytorchFI [27] and TensorFI [32, 33], enabling FI simulations in PyTorch and TensorFlow respectively. These software-level FI tools support various fault models and provide different metrics for the resilience analysis of DNNs regarding parameters and activations.
Any FI simulation requires multiple simulations while a random set of parameters/activations are faulty and propagated through the forward execution of a DNN. Then, the erroneous outputs are observed, and reliability metrics are acquired. The number of simulations affects the accuracy of the obtained metrics. In exhaustive FI, all possible bitflips in the fault space are flipped individually, leading to a huge simulation space, which is impractical since DNNs possess millions of parameters and activations. Statistical Fault Injection (SFI) attempts to reduce the number of simulations based on statistical analysis, while guaranteeing a certain maximum error margin.
For conventional computing hardware, the number of simulations in an SFI experiment used to be specified by Eq. (1), where n is the number of samples, N is total fault space, e is the error and t is determined based on the expected confidence-level [34].
@@
However, it is shown that applying Eq. (1) to the entire DNN does not result in statistically accurate results [26]. Authors in [26] introduced multiple methods to improve the accuracy of SFI as well as to reduce the number of simulations. Accordingly, layer-wise SFI, data-unaware SFI and data-aware SFI were presented, which generally apply Eq. (1) to each layer separately. Data-unaware SFI and data-aware SFI methods consider FI at the bit level to improve the statistical experiments. In data-unaware p = 0.5 for all bits in Eq. (1), whereas in data-aware SFI requires a pre-analysis to specify the p for each bit position in the data representation.
@@
Yet, SFI requires hundreds of thousands of simulations in any FI method leading to a significantly long simulation time even with leveraging GPUs. Furthermore, the growth of emerging DNNs necessitates an increasing number of simulations. Therefore, reliability analysis requires a paradigm shift in resilience analysis. Analytical methods for reliability analysis of DNNs are proposed to address its scalability issue.
@@
Throughout the literature, multiple techniques are proposed to obtain the resilience of CNNs based on non-FI approaches, called analytical methods [22]. Layer-wise Relevance Propagation (LRP) is proposed to derive the contribution of neurons to the classification output of CNNs expressing their criticality in the presence of faults and errors [35]. The gradient-based analysis is exploited in [36, 37, 38, 39, 40] to obtain the resilience of CNNs against faults. Nonetheless, gradients neither represent the impact of faults on misclassification nor result in a probabilistic outcome of faults’ effect on outputs. Also, faults may change the values significantly, while gradients represent small variations in erroneous values.
@@
Therefore, the existing analytical approaches do not result in accurate metrics for the reliability analysis of DNNs. DeepVigor was proposed to address this gap in the literature to provide fast and accurate metrics for the resilience of CNNs against faults [23]. It provides a Vulnerability Factor (VF) for bits, neurons and layers in CNNs by acquiring vulnerability values for each neuron (i.e., output feature maps). Vulnerability values represent how much a fault should change the golden value of a neuron’s output to misclassify CNN’s golden classification. Although DeepVigor is demonstrated to be faster than FI due to its fault propagation analysis method, it requires a high execution time to obtain VFs. The reason is that DeepVigor attempts to analyze all neurons in a CNN as well as to find the vulnerability values in a huge space of numbers.
This paper proposes DeepVigor+, a new method to quickly obtain VF metrics for layers and DNN models addressing both scalability and accuracy of fault resilience analysis in the literature. DeepVigor+ employs an optimized error propagation analysis in neurons, assuming that a single fault might happen either at its inputs or weights, thus leading to effectively shrinking the search space for vulnerability values. Moreover, DeepVigor+ proposes a stratified sampling to further accelerate the process of resilience analysis without a tangible analysis accuracy mitigation, leading to obtaining VF in a few minutes, even for deep and large emerging DNNs.

### v2

# DeepVigor+: Scalable and Accurate Semi-Analytical Fault Resilience Analysis for Convolutional Neural Networks

@@ Abstract
The growing exploitation of Machine Learning (ML) in safety-critical applications necessitates rigorous safety analysis. Hardware reliability assessment is a major concern with respect to measuring the level of safety in ML-based systems. Quantifying the reliability of emerging ML models, including Convolutional Neural Networks (CNNs), is highly complex due to their enormous size in terms of the number of parameters and computations. Conventionally, Fault Injection (FI) is applied to perform a reliability measurement. However, performing FI on modern-day CNNs is prohibitively time-consuming if an acceptable confidence level is to be achieved. To speed up FI for large CNNs, statistical FI (SFI) has been proposed, but its runtimes are still considerably long.
In this work, we introduce DeepVigor+, a scalable, fast, and accurate semi-analytical method as an efficient alternative for reliability measurement in CNNs. DeepVigor+ implements a fault propagation analysis model and attempts to acquire Vulnerability Factors (VFs) as reliability metrics in an optimal way. The results indicate that DeepVigor+ obtains VFs for CNN models with an error less than $1\%$, i.e., the objective in SFI, but with $14.9$ up to $26.9$ times fewer simulations than the best-known state-of-the-art SFI. DeepVigor+ enables an accurate reliability analysis for large and deep CNNs within a few minutes, rather than achieving the same results in days or weeks.
@@ I Introduction
In recent years, the rapid advancement of Artificial Intelligence (AI), in particular of Machine Learning (ML), has provided a unique opportunity for automation, leading to AI exploitation in safety-critical applications [1, 2]. Deep Neural Networks (DNNs) are dominantly employed in ML for complex tasks such as image classification and object detection that have applications in the safety-critical domain [3, 4]. According to recent regulations such as the European Union’s AI ACT [5] and the US National AI Initiative Act [6], AI systems that may cause adverse impacts on people’s safety are considered high-risk and must be rigorously assessed before deployment. This reflects the importance of early-stage safety risk assessment for AI deployment in safety-critical applications.
One of the primary concerns in safety-critical applications is ensuring the hardware reliability of the underlying computational platforms concerning random hardware faults [7]. In the context of DNN accelerators, hardware reliability refers to the probability that a DNN accelerator performs as intended during the deployment, despite potential faults in hardware [8]. Reliability assessment typically assumes the DL model and input data are error-free, focusing mainly on the impact of hardware faults on the application [9]. Resilience analysis is the process of studying the impact of faults on the application’s results [10]. Such faults appear due to soft errors, electromagnetic, temperature variation, aging, etc., that can perturb a DNN’s parameters or activation values during runtime, potentially leading to catastrophic consequences. As an example, a hardware fault may result in failing the detection of objects on the road and result in a fatal accident, as shown in Fig. 1.
With the technology miniaturization, the rate at which hardware faults occur is continuously rising in logic and memory circuits since the sensitivity of transistors increases against soft errors, temperature variations, etc. [11, 12, 13]. This trend significantly impacts the reliability of DNN accelerators, especially as they are increasingly adopted in applications such as perception modules in autonomous vehicles [14]. Therefore, assessing the hardware reliability of DNNs during the design and development phases is necessary, not only to meet safety certification requirements but also to proactively address potential reliability issues before deployment.
Figure 1: A real-world example of an autonomous vehicle failure, leading to a fatal accident, source: https://shorturl.at/QnOQq.
@@
On the other hand, the size of DNNs in terms of memory footprint and number of operations is rapidly growing to achieve higher accuracy [16, 17]. Fig. 2 presents the growing model size in terms of the number of MAC operations and parameters for emerging Convolutional Neural Networks (CNNs) and Vision Transformers (ViT) [15]. This trend necessitates leveraging more complex and performant computing systems such as GPUs, TPUs, etc. [18]. The fact that these complex systems are prone to hardware faults [19, 20, 21] leads to a necessary and challenging task of their hardware reliability assessment prior to deployment in safety-critical applications.
@@
To ensure a reliable DNN deployment, the first step is to extensively evaluate the performance and functionality of pre-trained DNNs against hardware faults. In this regard, several research works have been carried out to assess and enhance DNNs’ hardware reliability with respect to hardware faults listed in the related surveys [19, 22, 23, 24, 10]. To achieve this objective, three main approaches are identified throughout the state-of-the-art, in a prior systematic literature review [10]: 1) Fault Injection (FI) methods in which faults are injected into the target system and simulated, 2) Analytical methods where faults effects are mathematically analyzed, and 3) Hybrid methods that combine FI and analytical methods.
@@
It was observed that the majority of existing research works adopt FI (based on simulation, emulation, or irradiation). In FI experiments, faults can be accurately modeled, and the behavior of a system is evaluated in the presence of faults. However, given the growing size of emerging DNNs and their accelerators [18, 25], obtaining a fully precise evaluation by FI requires billions of simulations, which takes weeks of execution, making it unfeasible and impractical.
@@
Throughout the literature, multiple research works are presented to significantly reduce the simulation complexity of FI experiments for DNNs while maintaining statistical accuracy. It includes software simulation [26, 27, 28, 29], or hardware emulation [30, 31, 32]. Nonetheless, any FI experiment requires a certain level of statistical confidence, which is obtained by a considerable number of repetitions and a multitude of fault locations [27]. Therefore, FI-based approaches for reliability assessment are inherently non-scalable and they take days and weeks using powerful computing resources such as GPUs.
On the other hand, analytical approaches are presented to tackle the scalability issue. Nonetheless, they cannot provide reliability-oriented metrics, including Vulnerability Factors (VFs), which are an essential metric for reliability evaluation. VF is the probability that a fault in a particular structure will result in an error. VF is in line with the probabilistic nature of reliability evaluation [33], and it is widely used for reliability analysis of DNNs [10]. However, existing analytical methods attempt to achieve a relative contribution of neurons/weights to construct the classification outputs, and do not represent the impact of faults on the outputs of DNNs in a probabilistic way. Therefore, the accuracy of analytical methods is not comparable to that of FI-based methods.
@@
To tackle the existing issues in the literature, a prior semi-analytical method named DeepVigor [8] is proposed as an alternative to simulation-based FI. DeepVigor employs analytical fault propagation combined with simulation-based computations, categorized as ”semi-analytical”, providing accurate VF for bits, neurons, and layers of CNNs. It is demonstrated that DeepVigor is faster than FI by utilizing a fault propagation analytical method. However, it analyzes all neurons in a CNN as well as exploits a huge search space of real numbers to obtain vulnerability values, leading to a scalability issue for large and deep CNNs.
@@
In this paper, we propose a scalable, fast, and accurate alternative to FI named DeepVigor+, which provides VF for layers and models of CNNs. DeepVigor+ takes advantage of a theoretical fault propagation analysis within neurons and entire DNNs to derive VFs in an optimized way. To the best of our knowledge, DeepVigor+ is the first scalable semi-analytical alternative to FI with a comparable accuracy for resilience analysis of DNNs in the literature. The contributions of this paper are as follows:
@@
  • Introducing a scalable, fast, and accurate resilience analysis method for CNNs called DeepVigor+ for deriving VFs for convolutional layers by analyzing both parameters and activations. DeepVigor+ is built on a theoretical ground to enable an optimal error propagation analysis. Moreover, it conducts a novel statistical approach based on stratified sampling, unleashing a fast and accurate resilience analysis for large and deep CNNs,
@@
  • Extensively demonstrating the effectiveness of DeepVigor+ by examining its error compared to FI, leading to meeting a $1\%$ mean absolute error in obtaining VF. Furthermore, its scalability is evidenced against statistical FI, resulting in $14.9$ to $26.9$ times fewer simulations than the best state-of-the-art statistical FI approach (i.e., data-aware statistical FI [27]). Derived VF by DeepVigor+ can provide the possibility of reliability comparison and visualization between and within various DNN models in a few minutes for deep and large DNNs.
@@
DeepVigor+ is presented as an open-source tool (https://github.com/mhahmadilivany/DeepVigor), unlocking reliability analysis for emerging DNNs and enabling researchers to quickly assess their reliability and develop fault-tolerant DNNs.
@@
Table I presents the abbreviations frequently used in the manuscript. In the rest of the paper, Section II highlights the gap by overviewing the related works on statistical FI as well as analytical methods. Section III presents the methodology of DeepVigor+. Section IV explains the implementation of DeepVigor+ and Section V presents the results. Section VI presents the limitations and sheds light on the potential impacts of DeepVigor+ for reliability analysis of emerging DL systems, and finally, Section VII concludes the paper.
@@ II Related Works: Resilience Analysis Methods for DNNs
With the fast growth of emerging DNNs’ size regarding their memory and computation requirements, their reliability analysis becomes prohibitively complex, time-consuming, and resource-hungry. In a recent literature review [10], it has been observed that the majority of papers studying the reliability of DNNs exploit a Fault Injection (FI) approach. Software-level simulation FI is widely adopted in reliability analysis due to its low design time and fast execution.
Multiple open-source tools for software-level simulation-based reliability analysis are proposed in the literature, and the most adopted ones include PytorchFI [28] and TensorFI [34, 35], enabling FI simulations in PyTorch and TensorFlow, respectively. These software-level FI tools support various fault models and provide different metrics for the resilience analysis of DNNs regarding parameters and activations.
Any FI experiment requires multiple simulations while a random set of parameters/activations are being faulty and propagated through the forward execution of a DNN. Then, the erroneous outputs are observed, and reliability metrics are acquired. The number of simulations affects the accuracy of the obtained metrics. In exhaustive FI, all the bits in the fault space are to be flipped and simulated individually, leading to a huge simulation space, which is impractical since DNNs possess millions of parameters and activations. Statistical Fault Injection (SFI) attempts to reduce the number of simulations based on statistical analysis, while guaranteeing a certain maximum error margin.
For conventional computing hardware, the number of simulations in an SFI experiment is specified by Eq. (1), where n is the number of samples, N is the total fault space, e is the error, and t is determined based on the expected confidence-level [36].
@@
However, it is shown that applying Eq. (1) to the entire DNN does not result in statistically accurate results [27]. Authors in [27] introduce multiple methods to improve the accuracy of SFI as well as to reduce the number of simulations. Accordingly, layer-wise SFI, data-unaware SFI, and data-aware SFI are presented, which generally apply Eq. (1) to each layer separately. Data-unaware SFI and data-aware SFI methods consider FI at the bit level to improve the statistical experiments. In data-unaware p = 0.5 for all bits in Eq. (1), whereas in data-aware SFI requires a pre-analysis to specify the p for each bit position in the data representation.
@@
Yet, SFI requires hundreds of thousands of simulations, leading to a significantly long simulation time, in the order of weeks, even with leveraging GPUs. Furthermore, the growth of emerging DNNs necessitates an increasing number of simulations. Therefore, reliability analysis requires a paradigm shift in resilience analysis. Analytical methods for reliability analysis of DNNs are proposed to address the scalability issue.
@@
Throughout the literature, multiple techniques are proposed to obtain the resilience of CNNs based on non-FI approaches, called analytical methods [10]. Layer-wise Relevance Propagation (LRP) is proposed to derive the contribution of neurons to the classification output of CNNs, expressing their relative criticality in the presence of faults and errors [37]. However, this notion of criticality primarily captures the potential impact of a neuron’s malfunction on the output, rather than providing a probabilistic measure of classification errors due to faults. As such, it does not directly reflect the neuron’s reliability.
The gradient-based analysis is exploited in [38, 39, 40, 41, 42] to obtain the resilience of CNNs against faults. Nonetheless, gradients do not represent the impact of faults on classification and do not result in a probabilistic outcome of the faults’ effect on the outputs. Also, faults may change the values significantly, while gradients represent small variations in erroneous values.
@@
Thus, the existing analytical approaches do not result in accurate metrics for the reliability analysis of DNNs. DeepVigor was proposed to address this gap in the literature to provide fast and accurate metrics for the resilience of CNNs against faults [8]. It provides a Vulnerability Factor (VF) for bits, neurons, and layers in CNNs by acquiring vulnerability values for each neuron (i.e., output feature maps). Vulnerability values represent how much a fault should change the golden value of a neuron’s output to change the CNN’s golden classification.
Although DeepVigor obtains VFs with a high accuracy (i.e., $0.1\%$) faster than FI, there are a few drawbacks in its method, which make it non-scalable for deep and large CNNs. The drawbacks of DeepVigor and the ways DeepVigor+ addresses them are as follows:
  • To obtain VFs, DeepVigor first derives vulnerability value ranges with a huge search space based on the real numbers. However, 32-bit floating-point values do not represent all possible real numbers. To address this, DeepVigor+ builds a theoretical ground for optimal error propagation and then constructs a vulnerability value search space that includes all possible errors produced due to bitflips in 32-bit floating-point data representation.
  • To derive VFs, DeepVigor performs a bitflip mapping step in its operation, which conducts an exhaustive bitflip for each neuron to map the produced errors to vulnerability value ranges. This step is another reason that hinders DeepVigor from scaling. To tackle it, in this paper, we propose a mathematical approach to compute VFs from vulnerability value ranges with minimal computations based on the distribution of errors.
  • DeepVigor analyzes all neurons in a CNN, i.e., Output Feature Maps (OFMaps), in layers. This exhaustive analysis hinders a fast resilience analysis for large and deep CNNs and makes it impractical. To address this, we propose a stratified sampling method in DeepVigor+ to address the scalability of resilience analysis, while meeting the statistical requirements (i.e., $1\%$).
  • DeepVigor only analyzes the activations to achieve VF and excludes analyzing the fault resilience of weights. To address this gap, we analyze both activations and weights in DeepVigor+ and obtain the Model Vulnerability Factor (MVF) for an entire CNN.